Shift Summary for Eve shift, 2/Sept/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Thu, 2 Sep 1999 21:30:38 -0500

Shift Summary for Eve shift, 2/Sept/99

On Shift: Hogan Nguyen, Miguel Barrio

Hours of beam available at all intensities: 3.0
Hours of beam available at nominal intensity: 3.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 3.0

Accounting of lost time:
a few kfrends, but that accounts for most of the lost time on our
part...

========================================================================
Runs ended this shift:
Run 14408 - 30.4M events - BEAM_MAGOFF

Tapes brought to FCC this shift: QKE019-QKE027
QKB129-QKB130
QKS372-QKS374

========================================================================
Accesses during shift:
none

========================================================================
Overview of shift:
Special Run: hadron run with magnet off, B11 prescale = 2

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14408 1:45 Started to get beam for da special run,
at SEM = 6.8E12. After a couple of trials
to fine-tune the B11 prescale (to maximize
amount of data written to tape) we settled
for B11 prescale = 2.

Beam was on and off, for a total of about
three hours of data. MCR called a coupla times
reporting "main injector problems" and a
problem with a valve in the booster (these
where the reasons for lost beam time).

6:30 MCR calls to inform us that they will shut-down
early cuz they still have problems with the
main injector.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
Arun and Angela should read their email cuz there are a buncha things
to do during shutdown.